1. Suppose a university has an opportunity to build a solar facility just outside of the city that would generate about 120,000 MWh per year (1 MWh = 1000 KWh), and cost $25 million to install.

To determine if this is a good deal for the university, you need to analyze the economics. Assume the market price for electricity over the next five years is $48 per MWh and there is a 5-year project horizon. Evaluate this decision using a 3% discount rate. For this problem, assume that the costs of installing the solar panels occur at time 0 (beginning of year 1), and the benefits occur at the end of each subsequent year (end of year 1, end of year 2, etc.). The 5 year horizon implies that you only need to evaluate the gains of the solar project over a 5 year period. Solar panels obviously will provide benefits in the future if they remain installed, but you should ignore years beyond the 5th year for this analysis.

The present value of the stream of costs is $ _______ million and the present value of the stream of benefits is $ ________ million.

A. 0;28.8

B. 25; 23.5

C.25; 26.4

D. 25; 28.8

2. Use the data for the last question as you consider this question. Suppose Ohio State University sells the resulting renewable energy credits from this project on the market at the current price of renewable energy credits in Ohio of $5 per MWh. How much would Ohio State make each year selling renewable energy credits (RECs), and is this a benefit or a cost to the University?

3.
Answer #1

i. The cost to install a project is $25000000, so the stream of cost is $25000000.

The cash flows in this project are same across five years = $48*120000= $5760000

This is the periodic stream of benefit from the project for next 5 years.

To calculate the present value when cash flows occur at the end of the year (remember payments can be made at the beginning of the year as well, in that case the formula will change), the formula to be used is =

PV= C*((1-(1+r)^-n))/r)

C= Periodic cash flow

n= No. of periods/payments to be made

r= interest rate

PV= 28,288,771

So, PV of stream of benefit is $28,288,771.

So, answer is option D

ii) In this case we have to calculate the NPV of the project, if NPV is positive, the project is beneficial

NPV = (C)/( 1+r) t - Initial investment

C= Cash flows in the time period= 5*120000

r = Discount rate= 3%

t = time period= 5

NPV = -22,252,175.69. Since it is negative, the project is not beneficial, it is the cost to the university.